OpenStudy (anonymous):

The product of an initial investment, I, and the quantity of the sum of 1 and the annual interest rate, r, raised to the power of n, the number of years of the investment, is equal to M, the current amount of money in an investment account. If an initial investment of $1,997.00 is made to an account with an annual interest rate of 7%, what will be the value of M after 6 years? Round to the nearest cent. $2,835.74 $2,996.96 $1,292.04 $2,800.90

OpenStudy (shamil98):

Use the formula: \[\huge F = Pe^{r*t}\]

OpenStudy (shamil98):

F = future value P = current value e = (math constant) and 2.71 r = rate t = time (in years)

OpenStudy (shamil98):

P = 1997 e = 2.71 r = 0.07 t = 6 plug it in and evaluate it

OpenStudy (anonymous):

so do i multiply and add

OpenStudy (shamil98):

you just multiply, use the formula i gave...

OpenStudy (anonymous):

im not really getting either answer im using P and E but not getting either

OpenStudy (anonymous):

i got over 5000

OpenStudy (anonymous):

what is it that im doing wrong

OpenStudy (shamil98):

\[\huge F = 1997(2.71^{.4})\] do it from here...

OpenStudy (anonymous):

ok im multiplying 2.71 4 times but still not getting it its over each answer chose i guess im not comprehending right

OpenStudy (shamil98):

you're not suppose to multiply 2.71 to .4 .4 is an EXPONENT

OpenStudy (anonymous):

oh i see now ive gotten 2268.4

OpenStudy (shamil98):

No... that's wrong ...

OpenStudy (anonymous):

u know wht it not helping and especially not helping telling people "how in the world do i not see that it is an exponent" im sorry im not like u sheesh

OpenStudy (anonymous):

just nevermind ill figure it out

OpenStudy (anonymous):

\[\large \color {green}{1,997\times \left(1.07\right)^6}\] and a calculator

OpenStudy (anonymous):

there is no \(e\) in this